4. Write the equation of the line passing through the points (-6,1) and (-4,2)
Answer by Alan3354(69443)

Do it like this one.
==========================
Write the equation of the line passing through the points (0-4) and (-2,2).
----------
The simplest way:
Step 1, find the slope, m
===========
m = diffy/diffx = -6/2 = -3
----
Step 2:
y-y1 = m*(x-x1) where (x1,y1) is either point.
---
y+4 = -3*(x)
3x + y = -4